OpenStudy (anonymous):

A random sample of 100 undergraduate students at a university found that 78 of them had used the university library’s website to find resources for a class. What is the margin of error for the true proportion of all undergraduates who had used the library’s website to find resources for a class?

OpenStudy (amistre64):

margin of error suggests a confidence interval, but no width has been suggested

OpenStudy (amistre64):

if we wanted to be 99% confident that the true population proportion was within a given interval, we would use z=2.576 and use the standard error as a multiplier

OpenStudy (anonymous):

what would the equation I would use be???

OpenStudy (queelius):

Hint: use the central limit theorem, and use a maximum value for the standard deviation for a bernouli random variable.

OpenStudy (queelius):

Now, as amistre indicated, we don't know how "confident" you want the estimate to be, but you can leave it unspecified, e.g., the alpha percentile.

OpenStudy (amistre64):

the standard error is just a modification of the standard deviation. in this case: and i dont really understand why, the binomial standard deviation is sqrt(npq) with a mean of np (the mean i get) \[z=\frac{x-np}{\sqrt{npq}}\] \[z=\frac{\frac{x-np}{n}}{\frac{\sqrt{npq}}{n}}\] \[z=\frac{\frac{x}{n}-p}{ \sqrt{\frac{npq}{n^2}}}\] \[z=\frac{\frac{x}{n}-p}{ \sqrt{\frac{pq}{n}}}\] \[z \sqrt{\frac{pq}{n}}=\frac{x}{n}-p\] \[\underbrace{p}_{midpoint}\pm \underbrace{z \sqrt{\frac{pq}{n}}}_{margin~of~error}=\underbrace{\pm\frac{x}{n}}_{end~points}\]

OpenStudy (amistre64):

something like that,